OpenStudy (anonymous):

Ms. Smith invested $22,000 in two accounts, one yielding 5% interest and the other yielding 9%. If she received a total of $1740 in interest at the end of the year, how much did she invest in each account?

OpenStudy (mathstudent55):

I assume this is simple interest which is \(I = prt\) Here t = time = 1 year. Let x be the amount invested at 5% and y be the amount invested at 9%. The first equation represents the total amount invested: \(x + y = 22000 \) The second equation represents the interest earned: \(5%x + 9%y = 1740\), or \(0.05x + 0.09y = 1740\) Solve the system of equations for x and y: \(x + y = 22000 \) \(0.05x + 0.09y = 1740\)

OpenStudy (anonymous):

am i doing y=2200-x

OpenStudy (anonymous):

then .09y=.05x-2740?

OpenStudy (anonymous):

1740*

OpenStudy (mathstudent55):

Careful, \(y = 22000 - x\)

OpenStudy (mathstudent55):

Once you solve the first equation for y, then substitute what y is into the y of the second equation. \(x + y = 22000\) ----> \(y = 22000 - x\) \(0.05x + 0.09y = 1740\) \(0.05x + 0.09(22000 - x) = 1740 \) Now solve this equation for x.

OpenStudy (anonymous):

so i got x=6000

OpenStudy (anonymous):

then i plug x back into x+y=22000 ?

OpenStudy (mathstudent55):

Correct.

OpenStudy (anonymous):

x is 5% & y is 9% ?

OpenStudy (mathstudent55):

Look above. I wrote it down so we don't forget. It's in my first response.

OpenStudy (anonymous):

got it. thank you!